OpenStudy (anonymous):

Let a and b be integers, and let m be a positive integer. Then a ≡ b (mod m) if and only if a mod m = b mod m. Prove that:

hartnn (hartnn):

let \(a mod m = x\), \(b mod m = y\) also x =y , y-x= 0 so , a = pm +x similarly b = qm +y subtract these 2 equations b-a = (p-q)m + y-x let p-q =r b= mr +a giving \(a\equiv b~ mod~ m \)

hartnn (hartnn):

@ganeshie8 @mathmath333 check whether i did it right :P

hartnn (hartnn):

not sure whether its complete *if and only if*

ganeshie8 (ganeshie8):

that looks good to me for backward direction

ganeshie8 (ganeshie8):

\(\impliedby : \) \[a \pmod{m} = b\pmod {m} \implies a\equiv b \pmod{m}\] \(\implies : \) \[a\equiv b \pmod{m} \implies a \pmod{m} = b\pmod {m} \]

OpenStudy (mathmath333):

b-a = (p-q)m + y-x this should be b-a = (q-p)m + y-x

hartnn (hartnn):

oh right!

hartnn (hartnn):

forward dir. : we have a \(\equiv \) b mod m b = mr +a what next ?

OpenStudy (mathmath333):

^ that is the definition of \(b\equiv a(mod~ m)\)

hartnn (hartnn):

O.o when b is divided by m, the remainder is a so, a \(\equiv \) b mod m

hartnn (hartnn):

or did i get that wrong :P

OpenStudy (mathmath333):

well these are the theorms and its standard proof u will find in number theory books

OpenStudy (mathmath333):

and that proofs are long i think

hartnn (hartnn):

i just want to know a ≡ b mod m gives b = mr + a or not ...

OpenStudy (mathmath333):

\(14\equiv 4 (mod 10)\\~\\ 4=-10\times 1+14\)

OpenStudy (anonymous):

\(\rlap{\color{red}{\huge\bigstar}\huge \color{green}{ \text{Welcome to Open Study! }}\color{red}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{Welcome to Open Study! }}\color{yellow}\bigstar}\) @parkash